What is the sum of the arithmetic series below?
2 + 5 + 8 + ... + 56

Answer by stanbon(75887)

What is the sum of the arithmetic series below?
2 + 5 + 8 + ... + 56
-----
1st: What is the rank of the final term,i.e. 56?
Note: a(1) = 2 and d = 3
---
56 = 2 + (n-1)3
54 = (n-1)3
n-1 = 18
n = 19, so 56 is the 19th term.
------
2nd: Find the sum:
Sum = (19/2)(2+56) = 19(29) = 551
====================
Cheers,
Stan H.
====================